# From Text to Talk: Audio-Language Model Needs Non-Autoregressive Joint Training

### v3

@@ 3.4 Modality-Aware Attention Mechanism
Our attention design enforces a step-wise pattern across three content types: (1) the input prompt uses standard causal attention; (2) text tokens $\mathcal{T}_{m}$ apply strict causal attention to the prompt, all prior spans, and preceding tokens within their span; and (3) audio tokens $\mathcal{A}_{m}$ use hybrid attention—bidirectional within their own span and causal attention to the prompt and all earlier spans. This pattern enables parallel audio span training in a single forward pass while preventing cross-span interference. See Appendix A.5 for an illustration.
@@ 4.1 Experimental Setups
To effectively train and evaluate our proposed TtT framework, we follow prior works (Zeng et al., 2025; Ding et al., 2025; Long et al., 2025) and adopt a diverse collection of multi-task datasets, including ASR, TTS, audio chat, text chat, Automated Audio Captioning (AAC), Speech Emotion Classification (SEC), Acoustic Scene Classification (ASC), and interleaved text–audio data. In total, the combined training corpus contains approximately 6.3 million samples across these tasks. For evaluation, we focus on three representative capabilities: (1) conversational reasoning via Audio-QA, (2) cross-modal alignment via ASR, and (3) audio comprehension via AAC. To further validate end-to-end speech-to-speech performance in realistic conversational scenarios, we additionally evaluate on URO-Bench (Yan et al., 2025), a comprehensive benchmark that integrates reasoning, understanding, and oral conversation tasks. Full details on training data composition and evaluation protocol are provided in Appendix A.6 and Appendix A.7.
@@
To effectively evaluate our model on these tasks, we carefully design the evaluation protocol and metrics: (1) For Audio-QA, we introduce an ASR-LLM pipeline that transcribes the model’s spoken responses using language-specific ASR systems (Paraformer-zh for Chinese and Whisper-Large-v3 for English) and leverages a powerful LLM-as-a-Judge (Qwen3-235B-A30B) to assess semantic correctness against ground-truth answers; (2) For the ASR task, we directly measure transcription accuracy using Word Error Rate (WER); (3) For the AAC task, we adopt the evaluation prompt from CLAIR-A (Wu et al., 2024), using thinking model Qwen3-30B-A3B to judge the caption quality; (4) For URO-Bench, we directly use the official evaluation code and protocol to ensure a fair comparison with existing systems; More details are in Appendix A.7.1.
@@
We compare TtT with state-of-the-art audio-language models, including Moshi Défossez et al. (2024), SpeechGPT Zhang et al. (2023), Kimi-Audio Ding et al. (2025), VITA-Audio Long et al. (2025), LLaMA-Omni Fang et al. (2025), GLM-4-Voice Zeng et al. (2024), Mini-Omni Xie and Wu (2024b) and SLAM-Omni Chen et al. (2025) (detailed descriptions in Appendix A.8).
Model Configuration
We adopt the Qwen2.5-Base model as the backbone, experimenting with parameter scales of 1.5B and 3B, and fine-tune all parameters during training. For the audio components, we directly follow the audio tokenizer and decoder design introduced in GLM-4-Voice (Zeng et al., 2024). These modules have been shown to provide efficient and high-quality speech tokenization and synthesis, and they allow our framework to leverage strong audio modeling without requiring additional architectural modifications. The training details are provide in Appendix A.9.
4.2 Validating the Hybrid AR-NAR Architecture
Table 1: Comprehensive evaluation of TtT framework. Higher ($\uparrow$) is better for Audio-QA, lower ($\downarrow$) is better for ASR. Datasets abbreviations are available in Table 7
@@ 4.2 Validating the Hybrid AR-NAR Architecture
Table 1 (Main Results) provides the comparative results for the Audio-QA and ASR tasks. Our proposed TtT framework consistently outperforms both pure AR and NAR variants across all metrics. Specifically, at the 3B scale, TtT-3B surpassing Qwen2.5-3B (AR) by +3.04, +24.68, +5.93, and +10.91. For ASR tasks, TtT-3B yielding improvements of 42.41 and 58.36 absolute WER points over Qwen2.5-3B (AR). These substantial gains stem from our hybrid AR–NAR design: the NAR diffusion component enables efficient parallel denoising for tighter audio–text alignment, capturing audio’s inherent source-target dependencies, while AR text generation maintains coherent cross-modal conditioning and respects target-target dependencies. In contrast, purely NAR models perform notably worse due to order confusion from applying order-agnostic objectives to inherently sequential text-audio sequences. We also observe consistent scaling trends, where TtT-3B substantially outperforms TtT-1.5B across all tasks.
Table 2: Performance comparison on Audio-QA, ASR, and AAC tasks. Higher ($\uparrow$) is better for Audio-QA and AAC; lower ($\downarrow$) is better for ASR. Datasets abbreviations are available in Table 7
@@
To better understand the contribution of each training strategy in our hybrid AR-NAR framework, we perform an ablation study based on the full model TtT-3B (AR-NAR). The variant w/o BANOM corresponds to removing batchwise AR & NAR objective mixing from the full model, w/o PPM removes prefix preservation masking, and w/o SST removes stochastic span truncation. Table 1 ablation study part presents the detailed results of our ablation experiments. From these results we draw the following conclusions: (1) All three training strategies have a positive impact on model performance, and removing any one of them leads to clear degradation. For instance, on the LLaMAQuestions dataset, removing SST reduces the score from 34.68 to 10.20. This drop occurs because stochastic truncation mitigates positional bias in ⟨EOA⟩ prediction, forcing span termination by semantic content rather than position. Removing it weakens variable-length audio generation and reduces flexibility in conversational outputs. (2) Removing BANOM yields the largest performance degradation. For example, on the AISHELL-2 dataset, the performance decreases from 12.53 to 18.58 when the strategy is removed. This mechanism is essential for exposing text tokens to clean audio prefixes during training, better matching inference. Without it, the model faces sharper train–test discrepancy, weakening cross-modal consistency and alignment.
Effect of multimodal alignment pretraining.
To further investigate the effectiveness of our method on top of a multimodally aligned pretrained model, we perform large-scale multimodal pretraining based on the Qwen2.5-3B-Base model. Specifically, we construct a corpus of approximately 200B tokens covering ASR, TTS, text-only data, and interleaved text–audio data. The model is trained with a standard AR objective using a global batch size of 256 for 140k steps. This pretraining stage equips the backbone model (Qwen2.5-3B-Base) with strong cross-modal alignment ability before applying our hybrid AR–NAR learning framework. Table 1 compares the AR-only and AR–NAR frameworks under two different training strategies, specifically training directly from Qwen2.5-3B-Base without multimodal pretraining (TtT-3B) and initialization from the multimodally aligned pretrained model (Pretrain+AR and Pretrain+TtT). From the table, we observe that: (1) When trained directly from Qwen2.5-3B-Base, our TtT framework achieves comparable or even superior performance to the AR-only baseline, indicating that the hybrid AR–NAR design is already competitive without pretraining; (2) when applied on top of the multimodally aligned pretrained model, Pretrain+TtT consistently matches or surpasses Pretrain+AR across both Audio-QA and ASR tasks. These results demonstrate that TtT not only performs strongly from scratch, but also benefits significantly when built upon large-scale multimodal alignment pretraining. Having validated the effectiveness of our hybrid architecture and the benefits of multimodal pretraining, we now compare our best model (Pretrain+TtT) against state-of-the-art audio-language models.
@@ 4.3 Benchmarking Against State-of-the-Art Models
Building on the demonstrated strengths of our hybrid AR–NAR architecture and multimodal pretraining, we now evaluate Pretrain+TtT against state-of-the-art audio-language models. Tables 2 and 3 group results by model scale, distinguishing efficient models from large ones. Notably, Moshi does not support ASR, and the official releases of LLaMA-Omni and SLAM-Omni lack ASR prompting, hence no ASR results are reported. GLM-4-Voice exhibits poor ASR performance due to the absence of task-specific system prompts. Mini-Omni and SpeechGPT exhibit poor generalization to Chinese ASR tasks, as they are trained solely on English speech. Among efficient models, Pretrain+TtT achieves state-of-the-art performance across Audio-QA, ASR, and AAC. It substantially outperforms 0.5B baselines such as Mini-Omni and SLAM-Omni on Audio-QA and ASR. While SLAM-Omni reports higher AAC scores (54.52 on Clotho, 50.46 on MACS), its official implementation relies on a separate 7B Vicuna model fine-tuned specifically for AAC. Notably, Pretrain+TtT also exceeds several 7B-scale models on some tasks: it outperforms SpeechGPT (7B) across all Audio-QA and ASR benchmarks, and surpasses Moshi (7B) on WQ. and AE. tasks. These results demonstrate that our hybrid AR–NAR design enables a compact 3B model to match or exceed the task-specific capabilities of significantly larger systems.
Beyond the standard benchmarks, we further validate our model on URO-Bench (Yan et al., 2025) a comprehensive speech-to-speech benchmark that assesses speech understanding, reasoning, and oral conversation across basic and pro difficulty levels. As shown in Table 3, among efficient models, Pretrain+TtT achieves the best performance across both basic and pro difficulty levels. Compared to large models, Pretrain+TtT outperforms Moshi (7B) and SpeechGPT (7B) across all task categories, and achieves comparable performance to VITA-Audio (7B) and LLaMA-Omni (8B) on pro reasoning tasks and pro oral conversation tasks. While GLM-4-Voice (9B) and Kimi-Audio (7B) achieve the highest scores overall, the performance gap is reasonable given their 3x model size. The perceptual quality of both TtT and Pretrain+TtT falls within the 3.89–4.5 range (NMOS & UTMOS), confirming consistently good audio synthesis quality. However, Kimi-Audio exhibits notably lower perceptual quality (UTMOS: 2.93, NMOS: 3.52) despite its strong task completion performance. This degradation stems from language consistency issues: Kimi-Audio frequently generates mixed Chinese-English audio or produces Chinese audio for English tasks. While the semantic content may be correct, such cross-lingual inconsistencies significantly degrade perceptual audio quality.
@@ 5 Conclusion
In this work, we introduce a unified framework that combines autoregressive text generation with non-autoregressive audio diffusion. By explicitly respecting the asymmetry between text and audio dependencies, our framework bridges the strengths of AR and NAR modeling within a single Transformer. We further propose simple yet effective strategies to mitigate train–test discrepancies, enabling robust and flexible audio generation. Experiments on Audio-QA and ASR benchmarks demonstrate clear improvements over strong AR and NAR baselines. Our results highlight the importance of modality-aware design for building scalable and effective speech-to-speech systems.
Reproducibility statement
The anonymous downloadable source code is available at: https://anonymous.4open.science/r/TtT. For theoretical results, a complete proof of the claim in included in the Section A.1.1 in Appendix.
@@ A.2.1 Audio-Language Model Pretraining
Recent advances in end-to-end audio-language models have moved beyond traditional cascaded architectures (Chen et al., 2022; Wang et al., 2023) toward unified multimodal frameworks. Representative works include Moshi (Défossez et al., 2024), which achieves real-time duplex speech conversation through hierarchical Transformer architectures; GLM4-Voice (Zeng et al., 2024), which builds upon GLM-4-9B for robust Chinese and English speech processing; and VITA-Audio (Long et al., 2025), which introduces a lightweight Multiple Cross-modal Token Prediction (MCTP) module for fast audio-text generation with significantly reduced first-token latency. More recent efforts have focused on scaling and production readiness: Step-Audio (Huang et al., 2025) presents a 130B-parameter unified speech-text model with generative speech data engine and instruction-driven fine control across dialects, emotions, singing, and RAP, while Baichuan-Audio (Li et al., 2025) features text-guided aligned speech generation with multi-codebook discretization to preserve both semantic and acoustic information. UniWav (Liu et al., 2025) proposes the first unified encoder-decoder framework that jointly learns representation encoders and generative audio decoders for both discriminative and generative speech tasks.
A key limitation shared by these approaches is their reliance on uniform autoregressive objectives for both text and audio tokens, which overlooks the distinct dependency structures of these modalities. Our work addresses this gap by proposing a hybrid AR-NAR framework that respects the inherent asymmetries between text and audio generation.
@@ A.2.2 Discrete Diffusion Models
Discrete diffusion models have emerged as a compelling alternative to autoregressive generation, offering non-autoregressive approaches that can generate entire sequences in parallel. The foundational work of D3PMs (Austin et al., 2021) generalized diffusion processes to discrete data through flexible transition matrices, with absorbing processes that progressively mask tokens proving particularly effective. This framework has since evolved through both theoretical advances and practical improvements. From a theoretical perspective, recent work has deepened our understanding of discrete diffusion dynamics. Ou et al. (2024) revealed that absorbing diffusion’s concrete score can be expressed as time-independent conditional probabilities, leading to RADD—a reparameterized model that removes explicit time conditioning while establishing connections to any-order autoregressive generation. Building on this foundation, Li and Cai (2025) formally characterized convergence rates, proving that KL divergence decays at $O(1/T)$ with bounds scaling linearly with token mutual information. However, Feng et al. (2025) identified a fundamental trade-off: while masked diffusion achieves near-optimal perplexity in constant steps, sequence-level tasks like reasoning may require steps linear in sequence length. Practical advances have focused on training efficiency and application domains. Shi et al. (2024) reformulated the variational objective as a weighted integral of cross-entropy losses, unifying prior approaches while achieving state-of-the-art results that even surpass comparable autoregressive baselines. For complex reasoning tasks where autoregressive models struggle with subgoal imbalance, Ye et al. (2024) demonstrated that Multi-Granularity Diffusion Modeling can achieve near-perfect accuracy by prioritizing harder subgoals during training. The scalability challenge has been addressed through innovative adaptation strategies. Rather than training from scratch, Gong et al. (2024); Nie et al. (2025) showed that pretrained autoregressive models can be efficiently converted to diffusion models via continual pre-training, maintaining competitive performance while enabling parallel generation. Meanwhile, hybrid approaches are gaining traction: Lovelace et al. (2024) combined diffusion-based latent proposals with autoregressive decoding for controllable generation, while Yang et al. (2025b) developed MMaDA, a unified multimodal diffusion foundation model that processes text, images, and reasoning within a single architecture.
A.3 Autoregressive Modeling & Absorbing Discrete Diffusion
A.3.1 Autoregressive Modeling
Autoregressive (AR) models are a fundamental class of generative models that factorize the joint probability distribution of a sequence $x=(x^{1},\ldots,x^{L})$ into a product of conditional probabilities, based on the chain rule:

### v4

@@ 3.4 Modality-Aware Attention Mechanism
Our attention design enforces a step-wise pattern across three content types: (1) the input prompt uses standard causal attention; (2) text tokens $\mathcal{T}_{m}$ apply strict causal attention to the prompt, all prior spans, and preceding tokens within their span; and (3) Audio tokens $\mathcal{A}_{m}$ use hybrid attention, attending bidirectionally within the span and causally to the prompt and all earlier spans. This enables parallel audio-span training in a single forward pass while preventing cross-span interference. See Appendix A.5 for an illustration.
@@ 4.1 Experimental Setups
To effectively train and evaluate our proposed TtT framework, we follow prior work (Zeng et al., 2025; Ding et al., 2025; Long et al., 2025) and adopt a diverse collection of multi-task datasets, including ASR, TTS, audio chat, text chat, AAC, speech emotion classification (SEC), acoustic scene classification (ASC), and interleaved text-audio data. In total, the combined training corpus contains approximately 6.3 million samples across these tasks. For evaluation, we focus on three representative capabilities: (1) conversational reasoning via Audio-QA, (2) cross-modal alignment via ASR, and (3) audio comprehension via AAC. To further validate end-to-end S2S performance in realistic conversational scenarios, we additionally evaluate on URO-Bench (Yan et al., 2025), a comprehensive benchmark that integrates reasoning, understanding, and oral conversation tasks. Full details on training data composition and evaluation protocol are provided in Appendix A.6 and Appendix A.7.
@@
To effectively evaluate our model on the above tasks, we carefully design the evaluation protocol and metrics: (1) For Audio-QA, we introduce an ASR-LLM pipeline that transcribes the model’s spoken responses using language-specific ASR systems (Paraformer-zh for Chinese and Whisper-Large-v3 for English) and leverages a powerful LLM-as-a-Judge model (Qwen3-30B-A3B [^1] https://huggingface.co/Qwen/Qwen3-30B-A3B-Thinking-2507) to assess semantic correctness against ground-truth answers; (2) For the ASR task, we directly measure transcription accuracy using word error rate (WER); (3) For the AAC task, we adopt the evaluation prompt from CLAIR-A (Wu et al., 2025), using thinking model Qwen3-30B-A3B to judge the caption quality; (4) For URO-Bench, we directly use the official evaluation code and protocol to ensure a fair comparison with existing systems. More details are in Appendix A.7.1.
@@
We compare TtT with state-of-the-art audio-language models, including Moshi Défossez et al. (2024), SpeechGPT Zhang et al. (2023), Kimi-Audio Ding et al. (2025), VITA-Audio Long et al. (2025), LLaMA-Omni Fang et al. (2025), GLM-4-Voice Zeng et al. (2024), Mini-Omni Xie and Wu (2024) and SLAM-Omni Chen et al. (2025a) (detailed descriptions in Appendix A.8).
Table 1: Comprehensive evaluation of TtT framework. Higher ($\uparrow$) is better for Audio-QA, lower ($\downarrow$) is better for ASR. Datasets abbreviations are available in Table 7.
@@
Model Configuration
We adopt the Qwen2.5-Base model as the backbone, experimenting with parameter scales of 1.5B and 3B, and fine-tune all parameters during training. For the audio components, we directly follow the audio tokenizer and decoder design introduced in GLM-4-Voice (Zeng et al., 2024). These modules have been shown to provide efficient and high-quality speech tokenization and synthesis, and they allow our framework to leverage strong audio modeling without requiring additional architectural modifications. The training details are provided in Appendix A.9.
4.2 Validating the Hybrid AR-NAR Architecture
@@
Table 1 (Main Results) provides the comparative results for the Audio-QA and ASR tasks. Our proposed TtT framework consistently outperforms both pure AR and NAR variants across all metrics. Specifically, at the 3B scale, TtT-3B surpasses Qwen2.5-3B (AR) by +3.04, +24.68, +5.93, and +10.91 on Audio-QA task. For ASR tasks, TtT-3B yields improvements of 42.41 and 58.36 absolute WER points over Qwen2.5-3B (AR). These substantial gains stem from our hybrid AR-NAR design: the NAR diffusion component enables efficient parallel denoising for tighter audio-text alignment, capturing audio’s inherent source-target dependencies, while AR text generation maintains coherent cross-modal conditioning and respects target-target dependencies. Purely NAR models perform notably worse due to order confusion from applying order-agnostic objectives to inherently sequential text-audio sequences.
Table 2: Performance comparison on Audio-QA, ASR, and AAC tasks. Higher ($\uparrow$) is better for Audio-QA and AAC; lower ($\downarrow$) is better for ASR. Datasets abbreviations are available in Table 7.
@@
To better understand the contribution of each training strategy in our hybrid AR-NAR framework, we perform an ablation study based on the full model TtT-3B (AR-NAR). The variant w/o BANOM corresponds to removing batchwise AR & NAR objective mixing from the full model, w/o PPM removes prefix preservation masking, and w/o SST removes stochastic span truncation. Table 1 (ablation study) presents the detailed results of our ablation experiments. From these results, we draw the following conclusions: All three training strategies have a positive impact on model performance, and removing any one of them leads to clear degradation on most metrics. For instance, on the LQ. dataset, removing SST reduces the score from 34.68 to 10.20. This drop occurs because stochastic truncation mitigates positional bias in ⟨EOA⟩ prediction, forcing span termination by semantic content rather than position. Removing it weakens variable-length audio generation and reduces flexibility in conversational outputs. Meanwhile, removing SST also substantially degrades ASR on major benchmarks, e.g., increasing WER on AISHELL-2 from 12.53 to 25.43 and on AISHELL-1 from 13.65 to 31.03.
Effect of Multimodal Alignment Pretraining
To further investigate the effectiveness of our method on top of a multimodally aligned pretrained model, we perform large-scale multimodal pretraining based on the Qwen2.5-3B-Base model. Specifically, we construct a corpus of approximately 200B tokens covering ASR, TTS, text-only data, and interleaved text-audio data. The model is trained with a standard AR objective using a global batch size of 256 for 140k steps. This pretraining stage equips the backbone model (Qwen2.5-3B-Base) with strong cross-modal alignment ability before applying our hybrid AR-NAR learning framework. Table 1 (training strategy comparison) compares the AR-only and AR-NAR frameworks under two different training strategies, specifically training directly from Qwen2.5-3B-Base without multimodal pretraining (TtT-3B) and initialization from the multimodally aligned pretrained model (Pretrain+AR and Pretrain+TtT). From the table, we observe that: (1) When trained directly from Qwen2.5-3B-Base, our TtT framework achieves comparable or even superior performance to the AR-only baseline, indicating that the hybrid AR-NAR design is already competitive without pretraining; (2) when applied on top of the multimodally aligned pretrained model, Pretrain+TtT consistently matches or surpasses Pretrain+AR across both Audio-QA and ASR tasks. These results demonstrate that TtT not only performs strongly from scratch, but also benefits significantly when built upon large-scale multimodal alignment pretraining.
@@ 4.3 Benchmarking Against State-of-the-Art Models
Having validated the effectiveness of our hybrid architecture and the benefits of multimodal pretraining, we now compare our best model (Pretrain+TtT) against state-of-the-art audio-language models. Tables 2 and 3 group results by model scale, distinguishing efficient models from large ones. Notably, Moshi does not support ASR, and the official releases of LLaMA-Omni, SLAM-Omni and GLM-4-Voice lack ASR prompting, hence no ASR results are reported. Mini-Omni and SpeechGPT exhibit poor generalization to Chinese ASR tasks, as they are trained solely on English speech. Among efficient models, Pretrain+TtT achieves state-of-the-art performance across Audio-QA and ASR, while remaining competitive on AAC. It substantially outperforms 0.5B baselines such as Mini-Omni on Audio-QA and ASR, and SLAM-Omni on Audio-QA. While SLAM-Omni reports higher AAC scores (54.52 on Clotho, 50.46 on MACS), its official implementation relies on a separate 7B Vicuna model fine-tuned specifically for AAC. Notably, Pretrain+TtT also exceeds several 7B-scale models on some tasks. It outperforms SpeechGPT on all ASR benchmarks and on AE. and LQ. tasks in Audio-QA, and surpasses Moshi on WQ. and AE. tasks. These results demonstrate that our hybrid AR-NAR design enables a compact 3B model to match or exceed some significantly larger systems on several benchmarks.
[…]
@@ 5 Conclusion
In this work, we introduce a unified framework that combines AR text generation with NAR audio diffusion. By explicitly respecting the asymmetry between text and audio dependencies, our framework bridges the strengths of AR and NAR modeling within a single Transformer. We further propose simple yet effective strategies to mitigate train-test discrepancies, enabling robust and flexible audio generation. Experiments on Audio-QA, ASR, AAC and URO-Bench benchmarks demonstrate clear improvements over strong AR and NAR baselines. Our results highlight the importance of modality-aware design for building scalable and effective S2S systems.
Acknowledgements
This work was supported in part by Beijing Natural Science Foundation under Grant No. L252010; in part by NFSC under Grant No. 62477025; in part by Key Laboratory of Smart Education of Guangdong Higher Education Institutes, Jinan University (2022LSYS003) and in part by the Excellent Graduate Student Cultivation Program of Jinan University under Grant No. 2024CXB040.
@@ A.2.1 Audio-Language Model Pretraining
Recent success of LLMs has been demonstrated through their exceptional performance in complex reasoning tasks, particularly in logical inference (Cheng et al., 2025; Liu et al., 2024) and mathematical problem-solving (Chen et al., 2025b; Liu et al., 2025b). Catalyzed by these advancements, there has been a paradigm shift towards general-purpose MLLMs capable of processing and generating information across diverse modalities (Zhu et al., 2025), along with broader considerations such as reliability and safety (Zhao et al., 2024; Li et al., 2025b; Zhao et al., 2025). Under this paradigm, recent advances in end-to-end audio-language models have moved beyond traditional cascaded architectures toward unified multimodal frameworks (Chen et al., 2022; Wang et al., 2023). Representative works include Moshi, which achieves real-time duplex speech conversation through hierarchical Transformer architectures (Défossez et al., 2024); GLM-4-Voice, which builds upon GLM-4-9B for robust Chinese and English speech processing (Zeng et al., 2024); and VITA-Audio, which introduces a lightweight multiple cross-modal token prediction (MCTP) module for fast audio-text generation with significantly reduced first-token latency (Long et al., 2025). More recent efforts have focused on scaling and production readiness: Step-Audio presents a 130B-parameter unified speech-text model with generative speech data engine and instruction-driven fine control across dialects, emotions, singing, and RAP (Huang et al., 2025), while Baichuan-Audio features text-guided aligned speech generation with multi-codebook discretization to preserve both semantic and acoustic information (Li et al., 2025a). UniWav proposes the first unified encoder-decoder framework that jointly learns representation encoders and generative audio decoders for both discriminative and generative speech tasks (Liu et al., 2025a).
A key limitation shared by these approaches is their reliance on uniform AR objectives for both text and audio tokens, which overlooks the distinct dependency structures of these modalities. Our work addresses this gap by proposing a hybrid AR-NAR framework that respects the inherent asymmetries between text and audio generation.
@@ A.2.2 Discrete Diffusion Models
Discrete diffusion models have emerged as a compelling alternative to AR generation, offering NAR approaches that can generate entire sequences in parallel (Chen et al., 2024). The foundational work of D3PMs generalized diffusion processes to discrete data through flexible transition matrices, with absorbing processes that progressively mask tokens proving particularly effective (Austin et al., 2021). This framework has since evolved through both theoretical advances and practical improvements. From a theoretical perspective, recent work has deepened our understanding of discrete diffusion dynamics. Ou et al. (2024) revealed that absorbing diffusion’s concrete score can be expressed as time-independent conditional probabilities, leading to RADD, which is a reparameterized model that removes explicit time conditioning while establishing connections to any-order AR generation. Building on this foundation, Li and Cai (2025) formally characterized convergence rates, proving that KL divergence decays at $O(1/T)$ with bounds scaling linearly with token mutual information. However, Feng et al. (2025) identified a fundamental trade-off: while masked diffusion achieves near-optimal perplexity in constant steps, sequence-level tasks like reasoning may require steps linear in sequence length. Practical advances have focused on training efficiency and application domains. Shi et al. (2024) reformulated the variational objective as a weighted integral of cross-entropy losses, unifying prior approaches while achieving state-of-the-art results that even surpass comparable AR baselines. For complex reasoning tasks where AR models struggle with subgoal imbalance, Ye et al. (2025) demonstrated that multi-granularity diffusion modeling can achieve near-perfect accuracy by prioritizing harder subgoals during training. The scalability challenge has been addressed through innovative adaptation strategies. Rather than training from scratch, Gong et al. (2025); Nie et al. (2025) showed that pretrained AR models can be efficiently converted to diffusion models via continual pre-training, maintaining competitive performance while enabling parallel generation. Meanwhile, hybrid approaches are gaining traction: Lovelace et al. (2024) combined diffusion-based latent proposals with AR decoding for controllable generation, while Yang et al. (2025b) developed MMaDA, a unified multimodal diffusion foundation model that processes text, images, and reasoning within a single architecture.
A.3 AR Modeling & Absorbing Discrete Diffusion
A.3.1 AR Modeling
AR models are a fundamental class of generative models that factorize the joint probability distribution of a sequence $x=(x^{1},\ldots,x^{L})$ into a product of conditional probabilities, based on the chain rule: